OpenStudy (carolinar7):

What is x^2+2x+1 factored

OpenStudy (carolinar7):

@andrearchat

OpenStudy (0_0youscareme):

this is history dude

OpenStudy (0_0youscareme):

not math

OpenStudy (carolinar7):

dont tell me the answer

OpenStudy (carolinar7):

@andrearchat

OpenStudy (carolinar7):

help me

OpenStudy (andrearchat):

2x + 1 + x2 = -1 + 1 Reorder the terms: 1 + 2x + x2 = -1 + 1 Combine like terms: -1 + 1 = 0 1 + 2x + x2 = 0 Factor a perfect square on the left side: (x + 1)(x + 1) = 0 Calculate the square root of the right side: 0 Break this problem into two subproblems by setting (x + 1) equal to 0 and 0.

OpenStudy (carolinar7):

I did not understand your explenation

OpenStudy (andrearchat):

maybe this will help http://www.algebra.com/algebra/homework/Radicals/Radicals.faq.question.116770.html

OpenStudy (carolinar7):

oh